Find the pH and % dissociation of HAc in a solution obtained by mixing 25ml of 0.2M HCl and 25ml of 0.2M NaAc. Given Ka for HAc is 1.78*10-5

Grade:12

1 Answers

vikas askiitian expert
13 years ago

  HCl               +                CH3COONA             =>               CH3COOH   +   NaCl

25*0.2 (millimoles)           25*0.2(mllimoles)      => 0                   0                                 (initial)

 0                                            0                                           25*0.2          25*0.2           (final)

total millimoles of acetic acid = 25*0.2

total volume = 50ml

molarity of acetic acid (C)= 25*0.2/50=0.1

degree of dissotiation = (Ka/c)1/2 = (1.78*10-4)1/2

                                                 =1.33%